# Glyph-ByT5: A Customized Text Encoder for Accurate Visual Text Rendering

### v1

@@ H. Typography Layout Planning with GPT-4
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption GPT4 Refer to caption Human Refer to caption GPT4 Refer to caption Human Refer to caption GPT4 Refer to caption Human
Figure 16: Illustrating the effect of font type blending.

### v2

@@ H. Typography Layout Planning with GPT-4
I. Visual Quality
To assess the visual quality, we report the FID scores of Glyph-SDXL on two benchmarks: the COCO benchmark, consisting of 5K text-image pairs, and the VisualParagraphy benchmark, consisting of 1K design-oriented text-image pairs that contain visual text in each image. The FID scores are $28.06$ and $47.96$, respectively. While our method’s FID score on COCO images is slightly higher than the $26.48$ achieved by SDXL, it is important to note that the FID metric inherently favors natural images over design images. This is due to the FID’s reliance on Inception-v3, which is pre-trained on the natural image dataset ImageNet.
J. Statistics over multiple runs
[…]
$\leq$20 chars $\leq$20-50 chars $\leq$50-100 chars $\geq$100 chars
Glyph-SDXL $93.20\pm 0.64$ $93.19\pm 0.67$ $91.23\pm 0.60$ $89.98\pm 0.51$
Table 15: Mean+std of word-level prediction over five runs.
[…]
$\leq$20 chars $\leq$20-50 chars $\leq$50-100 chars $\geq$100 chars
90% $79.76\pm 0.78$ $65.68\pm 1.07$ $68.64\pm 1.33$ $62.32\pm 1.65$
95% $79.76\pm 0.78$ $64.96\pm 1.32$ $50.40\pm 1.01$ $47.20\pm 1.41$
100% $79.76\pm 0.78$ $64.96\pm 1.32$ $45.04\pm 1.90$ $23.12\pm 1.67$
Table 16: Mean+std of success rate over five runs.